OpenStudy (elleblythe):

The variable x varies directly as the square of r and inversely as the cube of t. When r=3 and t=27, x=24. What is x when r is 2 and t is 8? Please show solution as much as possible

OpenStudy (anonymous):

so the equation would be \[y=\frac{ r ^{2}k }{ t ^{3}}\] so you want to find k to plug into the equation

OpenStudy (anonymous):

sorry replace y with x

OpenStudy (anonymous):

\[24=\frac{ 3^{2} k}{ 27^{3} }\] \[24=\frac{ 9 k}{ 19683 }\] 9k=472392 k=52488

OpenStudy (anonymous):

so you replace k into the equation \[x=\frac{ x^{2} 52488}{ t^{3} }\]

OpenStudy (anonymous):

then replace the other variables \[y=\frac{ 2^{2} 52488}{ 8^{3} }\] \[y=\frac{ 4(52488)}{512 }\] \[y=\frac{209952}{512}\] y=about 410

OpenStudy (elleblythe):

@insanitycat it's not an exact value?

OpenStudy (anonymous):

no its a long decimal

OpenStudy (anonymous):

if in fraction form it would be y=6561/16